Find "f(4)" if "f(x)=x^2-4x+3"


1
Expert's answer
2020-04-01T16:16:16-0400

"f(x)=x^2-4x+3"

We have to find f(4)

substituting 4 in the place of x in the above equation we get

"f(4)=4^2-4\cdot 4+3=16-16+3=3"